Identify the greater number, wherever possible, in each of the following?
4^3 or 3^4

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

The expansion of 4^3= 4 × 4 × 4 = 64
The expansion of 3^4= 3 × 3 × 3 × 3 = 81
Clearly,
64 < 81
So, 4^3< 3^4
Hence 3^4is the greater number.

Thanks